solve for subsitution
15x-10y=1
15y=2+5x
8x+5y+14
x=52-9y
solve equation by graphong
x+y=4
y-x=4
solve by elimination
8x-y=47
x+y=16

Answer by mananth(16949)

15x-10y=1.............1
15y=2+5x ---------2
/15
y=0.13+ 0.33x
Plug the value of y in (1)
15x-10(0.13+0.33x)=1
15x-1.33-3.33x=1
15x-3.33x=1+1.33
11.67x=2.33
/11.67
x=0.2
Plug the value of x in (1)
15x-10y =1
15*0.2-10y=1
3-10y=1
-10y=-2
/-10
y=0.2
.......................
8x+5y=14 .............1
9y=52-x ---------2
/9
y=5.78-0.11x
Plug the value of y in (1)
8x+5(5.78-0.11x)=14
8x+28.89-0.56x=14
8x-0.56x=14-28.89
7.44x=-14.89
/7.44
x=-2
Plug the value of x in (1)
8x+5y=14
8*-2+5y=14
-16+5y=14
5y=30
/5
y=6
..................................
x+y=4 .............1
-x+y =4 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-x-y=-4
-x+y=4
Add the two equations
-2x = 0
x=0
plug value of x in (1)
x+y=4
y=4
equation1
when x= 0 y=4(0,4)
when y =0,x=4(4,0)
Equation 2
when x= 0,y=4(0,4)
when y =0x=-4(-4,0)
...

................
8 x + -1 y = 47 .............1
1 x + 1 y = 16 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
8 x -1 y = 47
1 x 1 y = 16
Add the two equations
9 x = 63
/ 9
x = 7
plug value of x in (1)
8 x + -1 y = 47
56 + -1 y = 47
-1 y = 47 -56
-1 y = -9
y = 9